Assume all benefits accrue to the buyer and all costs are borne by the seller.

Prof. G. has 1 ticket to 4 different Raptors games. If he does not sell them, they will not be used by him or anyone else. (That is, his supply of Raptors tickets in perfectly inelastic at Q=4.)

Mirwan's demand is given by the following table:

Qd P
1 $18
2 $15
3 $10
4 $6

Zhuncheng's demand is given by the following table:

Qd P
1 $11
2 $7
3 $3
4 $1

If these tickets are efficiently allocated, how much surplus is created?
Answer #1

If tickets are efficiently allocated then 3 tickets will be sold to Mirwan and 1 ticket to Zhunchengs

total surplus=(18-4)+(15-4)+(10-4)+(11-4)=14+11+6+7=38

Thus surplus of 38 is created